the larger of two consecutive ingers is 7 greater than twice the smaller.
find the integers

Answer by checkley71(8403)

LET X & X+1 BE THE 2 INTEGERS.
X+1=2X+7
X-2X=7-1
-X=6
X=-6 ANSWER FOR TRHE SMALLER INTEGER.
-6+1=-5 FOR THE LARGER INTEGER.
PROOF:
-6+1=2*-6+7
-5=-12+7
-5=-5
-----------------------------------------------------------
ANSWER FOR THE FIRST PROBLEMIS:
(X+4)+2X=25
----------------------------------------------------------
SECOND PROBLEM:
N+(N-4)+4N=62